OpenStudy (anonymous):

At West Painting, they get about three calls a day asking for an estimate of the cost for having the interior of a house painted. To write up an estimate for the cost of a job, they need to know how much paint a job will take. If they average painting 3/4 of a room with 2/5 of a gallon of paint, then they can paint, on average, rooms per gallon.

OpenStudy (anonymous):

@EmoCookieMonster

OpenStudy (anonymous):

I'm not sure about this one, I'm sorry.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Maybe @MasterprojectHD ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

Let me know if you need help on anything else. x

OpenStudy (anonymous):

not sure if I understand the question correctly rooms per gallon = avg number of rooms / avg number of gallons avg number of rooms= 3/4 avg number of gallons = 2/5

OpenStudy (anonymous):

ok